# Investigating aerosols as a reconciliation mechanism for K2-18 b JWST MIRI and NIRISS/NIRSpec observations

### v1

@@
Investigating aerosols as a reconciliation mechanism for K2-18 b JWST MIRI and NIRISS/NIRSpec observations
@@ 2 Methods
We started by retrieving a flat spectrum to assess the statistical significance of any spectral features, for both dataset and combined dataset. Subsequently, we carried out retrievals including CH4, CO2, CO, H2O, NH3, C2H2, C2H4, H2CO, H2S, SO2, and HCN, which allowed us to identify the key absorbers for the combined analysis. We used absorption cross-sections computed by ExoMol (Yurchenko et al., 2011; Tennyson & Yurchenko, 2012; Barton et al., 2013; Yurchenko et al., 2014; Barton et al., 2014), specifically those from Chubb et al. (2021).
@@
The contribution of aerosols is accounted for using Mie scattering calculations. Previous strategies relied on the use of a simplified function to include aerosol scattering throughout the Rayleigh and Mie regimes (e.g., Kitzmann & Heng (2017)). This approach ignores the absorbing properties of the aerosols, encoded in the imaginary part of the complex refractive index, since absorption is usually negligible compared to scattering for small particles in the Visible and near-IR spectral range. In the present work, we include the absorbing properties of photochemical hazes that could explain the features observed in the K2-18 b observations at longer wavelengths in addition to the scattering effect at shorter wavelengths in the NIRISS SOSS/NIRSpec G395H range. The haze extinction is thus calculated using complex refractive index data obtained on laboratory haze analogues called tholins. The absorption properties of photochemical hazes are known to vary with the gas composition, temperature, pressure and irradiation efficiency (Brassé et al., 2015; Sciamma-O’Brien et al., 2023; Drant et al., 2024). We thus used 7 different datasets compiled from the original study of Khare et al. (1984) and the recent data of Drant et al. (2025) to explore the influence of the haze absorbing properties on the retrieval and final posterior distribution. We performed simulations with the refractive index data of exoplanet haze analogs that were produced using two different experimental setups (PAMPRE at LATMOS and COSmIC at NASA Ames) using only CH4 as a reactive gas. In addition, we performed simulations with the refractive index data obtained on Titan tholins using various experimental setups and N2/CH4 ratio. The different refractive index data are summarized in Table 1. In the simulations, the haze particles are assumed spherical with a lognormal distribution in size. The particle radius is therefore the only microphysical parameter included in the retrieval. The consideration of fractal particles would increase the number of free parameters, e.g. with the fractal dimension and number of spherical monomers, produce other possibilities to explain the haze extinction and thus generally increase degeneracy (e.g., Adams et al. (2019)).
@@ 3 Reconciling NIRISS SOSS/NIRSpec G395H and MIRI LRS observations with aerosols
Previous studies have retrieved high abundances of CH4, ranging from 1% to over 10% (Madhusudhan et al., 2023; Schmidt et al., 2025; Jaziri et al., 2025; Hu et al., 2025). Such significant methane levels suggest the formation of aerosols, such as Titan-like hazes via photochemistry (e.g., Trainer et al. (2006) and Arney et al. (2016)). Considering the modeled temperatures from previous studies (Charnay et al., 2021; Blain et al., 2021), along with our retrieved low temperatures, the formation of methane clouds is also plausible. These clouds may exhibit similar opacities to photochemical hazes generated in H2-dominated atmospheres. Both hazes and clouds could contribute to the flattening of the spectra at shorter wavelengths while preserving strong features at longer wavelengths. This could explain the differences in feature amplitudes observed between NIRISS SOSS/NIRSpec G395H and MIRI LRS observations.
@@
The log-evidence results in Table 2 indicate a strong preference for all models compared to a flat line. This is driven by the NIRISS SOSS/NIRSpec G395H data, which have already been confirmed to contain significant spectral features, most notably CH4. This is further supported by the retrieved metallicity, which remains consistent with values obtained using only NIRISS SOSS/NIRSpec G395H data (approximately 40 with a free-chemistry model).
We also observe that the combined analysis does not improve the CO2 detection, which is expected since CO2 has no noticeable feature in the mid-IR, and shows even less support for C2H4 compared to retrievals based solely on MIRI LRS data.
@@
Tholins exo1 contributes to the observed features across multiple spectral ranges: near 7m in the MIRI LRS range and around 3.4m in the NIRSpec G395H range (alongside CH4) as absorption dominates, and at shorter wavelengths in the NIRISS SOSS range, affecting the spectral slope as scattering dominates. Notably, the retrieved CH4 abundance decreases from 6.3×10-2 to 3.0×10-4 when considering these photochemical hazes compared to the reference model. The haze properties retrieved are also reasonable and in agreement with the modeling predictions of Lavvas et al. (2019) made for the temperature sub-Neptune GJ 1214 b. For the retrieval considering exo1 tholins, we constrain the particle radius in log unit to -1.65 (in m), the haze density in log unit to 11.22 (m-3), the mean pressure of the haze layer in log unit to -0.04 (in Pa), and the dP of the haze layer in log unit to 2.05 (in Pa), see Figure 7. The retrievals considering hazes also tend to increase the atmospheric temperature (see Figure 6) which along with the haze properties and the methane abundance affects the strength of the spectral features. Although this result reflects a degenerate behavior between different parameters influencing extinction, its physical meaning is reasonable as haze absorption in the UV-Visible tend to increase the atmospheric temperature of the upper atmosphere (e.g., Arney et al. (2016)). This increased temperature driven by the haze heating properties could inhibit the formation of methane clouds and prevent the co-existence with the solid organic hazes.
@@ Acknowledgements.
This project has received funding from the European Research Council (ERC) under the ERC OxyPlanets projects (grant agreement No. 101053033). This project acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 679030/WHIPLASH). T.D. acknowledges support by the 2024 NOMIS-ETH postdoc fellowship program. T.D. thanks the NOMIS foundation and ETH Zurich for funding via the 2024 Fellowship program.References
@@
- Arney et al. (2016) Arney, G. N., Domagal‑Goldman, S. D., Meadows, V. S., et al. 2016, Astrobiology, 16, 873
- Barton et al. (2014) Barton, E. J., Chiu, C., Golpayegani, S., et al. 2014, MNRAS, 442, 1821
- Barton et al. (2013) Barton, E. J., Yurchenko, S. N., & Tennyson, J. 2013, MNRAS, 434, 1469
- Bell et al. (2022) Bell, T., Ahrer, E.-M., Brande, J., et al. 2022, The Journal of Open Source Software, 7, 4503
@@
- Changeat et al. (2025) Changeat, Q., Bardet, D., Chubb, K., et al. 2025, arXiv e-prints, arXiv:2505.18715
- Charnay et al. (2021) Charnay, B., Blain, D., Bézard, B., et al. 2021, A&A, 646, A171
- Chubb et al. (2021) Chubb, K. L., Rocchetto, M., Yurchenko, S. N., et al. 2021, A&A, 646, A21
- Drant et al. (2024) Drant, T., Garcia‑Caurel, E., Perrin, Z., et al. 2024, A&A, 682, A6
- Drant et al. (2025) Drant, T., Sciamma-O’Brien, E., Jovanovic, L., et al. 2025, A&Ain review
- Gao et al. (2023) Gao, P., Piette, A. A. A., Steinrueck, M. E., et al. 2023, The Astrophysical Journal, 951, 96
@@
- Hu et al. (2025) Hu, R., Bello-Arufe, A., Tokadjian, A., et al. 2025, arXiv e-prints, arXiv:2507.12622
@@
- Khare et al. (1984) Khare, B. N., Sagan, C., Arakawa, E. T., et al. 1984, Icarus, 60, 127
@@
- Lavvas et al. (2019) Lavvas, P., Koskinen, T., Steinrueck, M. E., García Muñoz, A., & Showman, A. P. 2019, The Astrophysical Journal, 878, 118
- Luque et al. (2025) Luque, R., Piaulet-Ghorayeb, C., Radica, M., et al. 2025, arXiv e-prints, arXiv:2505.13407
- Madhusudhan et al. (2025) Madhusudhan, N., Constantinou, S., Holmberg, M., et al. 2025, ApJ, 983, L40
- Madhusudhan et al. (2023) Madhusudhan, N., Sarkar, S., Constantinou, S., et al. 2023, ApJ, 956, L13
- Schmidt et al. (2025) Schmidt, S. P., MacDonald, R. J., Tsai, S.-M., et al. 2025, arXiv e-prints, arXiv:2501.18477
- Sciamma-O’Brien et al. (2023) Sciamma-O’Brien, E., Roush, T. L., Rannou, P., Dubois, D., & Salama, F. 2023, The Planetary Science Journal, 4, 121
@@
- Tennyson & Yurchenko (2012) Tennyson, J. & Yurchenko, S. N. 2012, MNRAS, 425, 21
- Trainer et al. (2006) Trainer, M., Pavlov, A., DeWitt, H., et al. 2006, PNAS, 103, 18035
- Welbanks et al. (2025) Welbanks, L., Nixon, M. C., McGill, P., et al. 2025, arXiv e-prints, arXiv:2504.21788
- Wogan et al. (2024) Wogan, N. F., Batalha, N. E., Zahnle, K. J., et al. 2024, The Astrophysical Journal Letters, 963, L7
- Yurchenko et al. (2011) Yurchenko, S. N., Barber, R. J., & Tennyson, J. 2011, MNRAS, 413, 1828
- Yurchenko et al. (2014) Yurchenko, S. N., Tennyson, J., & Bailey, J. e. a. 2014, Proceedings of the National Academy of Science, 111, 9379
@@ Appendix A Dataset offset
All these offsets are fixed to reduce the parameter space and accelerate aerosol retrievals. Allowing them to vary would affect the precise values of the retrieved parameters but not the main conclusions regarding the impact of aerosols on the spectra.
@@ Appendix B Retrieval setup
The non-equilibrium chemistry retrieval was carried out using FRECKLL coupled with TauREx 3, without aerosols. Kzz has been set to 1010. Its value, within common range of values, has been shown in Jaziri et al. (2025) to not significantly affect the retrieval. The priors are detailed in Table 3. Due to its high computational cost and limited stability, we present only one non-equilibrium retrieval, applied to the MIRI LRS data, which contain relatively few data points. This allows us to assess the consistency of the MIRI LRS observations under non-equilibrium conditions.
@@ Appendix C Inconsistency between NIRISS SOSS/NIRSpec G395H and MIRI LRS observations
A retrieval using all molecules on the NIRISS SOSS/NIRSpec G395H data yields a metallicity of 40.56, demonstrating that, in our case, free-chemistry retrievals tend to bias metallicities toward lower values compared to non-equilibrium models. This bias arises from missing molecules, such as CO or N2, that may be abundant but have no prominent spectral features, and therefore cannot be well-constrained in the retrieval. Nevertheless, the MIRI LRS retrievals tend to favor lower metallicity to increase the scale height, allowing better fits to the larger observed features. This conclusion is valid because CH4 is the dominant gas.
These results are inconsistent with those from the NIRISS SOSS/NIRSpec G395H analysis and the non-equilibrium chemical model, both of which favor higher metallicity and a high CH4 abundance, as seen in Figure 3. In contrast, the MIRI LRS retrievals indicate low metallicity and do not show a preference for CH4. However, as shown in Table 4, none of the tested models are statistically favored over a flat line. At best, they fall within the ”weak preference” range. H2CO is unlikely to be dominant. Its retrieved abundance of 10-5 exceeds the 3 upper limit derived by Jaziri et al. (2025). C2H2 and C2H4 could plausibly be present, especially considering high C/O ratio and high metallicity. However, the retrievals favor C2H2 over CH4, which is inconsistent with the NIRISS SOSS/NIRSpec G395H data.

### v3

# Investigating aerosols as a way to reconcile K2-18 b JWST MIRI and NIRISS/NIRSpec observations

@@
Investigating aerosols as a way to reconcile K2-18 b JWST MIRI and NIRISS/NIRSpec observations
@@ 2 Methods
We started by retrieving a flat spectrum to assess the statistical significance of any spectral features, for both dataset and combined dataset. Subsequently, we carried out retrievals including CH4, CO2, CO, H2O, NH3, C2H2, C2H4, H2CO, H2S, SO2, and HCN, which allowed us to identify the key absorbers for the combined analysis. We included these molecules because they have consistent absorption cross-sections computed by ExoMol in hydrogen-rich atmospheres (Chubb et al., 2021). We acknowledge that our reduced sample of hydrocarbons may introduce some quantitative bias in the results. However, Welbanks et al. (2025) showed that many hydrocarbons contribute with similar significance, and that the most abundant species (C2H2, C2H4, and C2H6) likely have comparable abundances (see also Jaziri et al., 2025). For this reason, we prioritized consistency in the opacity data.
@@
The contribution of aerosols is accounted for using Mie scattering calculations. Previous strategies relied on the use of a simplified function to include aerosol scattering throughout the Rayleigh and Mie regimes (e.g., Kitzmann & Heng, 2017). This approach ignores the absorbing properties of the aerosols, encoded in the imaginary part of the complex refractive index, since absorption is usually negligible compared to scattering for small particles in the Visible and near-IR spectral range. In the present work, we include the absorbing properties of photochemical hazes that could explain the features observed in the K2-18 b observations at longer wavelengths in addition to the scattering effect at shorter wavelengths in the NIRISS SOSS/NIRSpec G395H range. The haze extinction is thus calculated using complex refractive index data obtained on laboratory haze analogues called tholins. The absorption properties of photochemical hazes are known to vary with the gas composition, temperature, pressure and irradiation efficiency (Brassé et al., 2015; Sciamma-O’Brien et al., 2023; Drant et al., 2024). We thus used 7 different datasets compiled from the original study of Khare et al. (1984) and the recent data of Drant et al. (2025) to explore the influence of the haze absorbing properties on the retrieval and final posterior distribution. We performed simulations with the refractive index data of exoplanet haze analogs that were produced using two different experimental setups (PAMPRE at LATMOS and COSmIC at NASA Ames) using only CH4 as a reactive gas. In addition, we performed simulations with the refractive index data obtained on Titan tholins using various experimental setups and N2/CH4 ratio. The different refractive index data are summarized in Table 1. In the simulations, haze particles are assumed spherical with a lognormal size distribution, making the particle radius the only microphysical parameter in the retrieval. Photochemical hazes can form fractal aggregates, which increase the effective radius and scattering efficiency. Accounting for aggregates would require additional poorly constrained parameters, such as the number of monomers per particle and the fractal dimension, increasing degeneracy in the retrievals (e.g., Adams et al., 2019). Therefore, only spherical particles were considered.
@@ 3 Reconciling NIRISS SOSS/NIRSpec G395H and MIRI LRS observations with aerosols
Previous studies have retrieved high abundances of CH4, ranging from 1% to over 10% (Madhusudhan et al., 2023; Schmidt et al., 2025; Jaziri et al., 2025; Hu et al., 2025). Such significant methane levels suggest the formation of aerosols, such as Titan-like hazes via photochemistry (e.g., Trainer et al., 2006 and Arney et al., 2016). Considering the modeled temperatures from previous studies (Charnay et al., 2021; Blain et al., 2021), along with our retrieved low temperatures, the formation of methane clouds is also plausible. These clouds may exhibit similar opacities to photochemical hazes generated in H2-dominated atmospheres. Both hazes and clouds could contribute to the flattening of the spectra at shorter wavelengths while preserving strong features at longer wavelengths. This could explain the differences in feature amplitudes observed between NIRISS SOSS/NIRSpec G395H and MIRI LRS observations.
@@
The combined analysis is driven by the NIRISS SOSS/NIRSpec G395H data, which have already been confirmed to contain significant spectral features, most notably CH4, contrary to the MIRI LRS data. This is further supported by the retrieved metallicity, which remains consistent with values obtained using only NIRISS SOSS/NIRSpec G395H data (approximately 40 with a free-chemistry model).
We also observe that the combined analysis does not improve the CO2 detection, which is expected since CO2 has no noticeable feature in the mid-IR, and shows even less support for C2H4 compared to retrievals based solely on MIRI LRS data. The retrieved abundance of CH4 is also consistent with previous data reduction analysis from Luque et al. (2025) (see Figure 8).
@@
Tholins exo1 contributes to the observed features across multiple spectral ranges: near 7m in the MIRI LRS range and around 3.4m in the NIRSpec G395H range (alongside CH4) as absorption dominates, and at shorter wavelengths in the NIRISS SOSS range, affecting the spectral slope as scattering dominates. Notably, the retrieved CH4 abundance decreases from 6.3×10-2 to 3.0×10-4 when considering these photochemical hazes compared to the reference model. The haze properties retrieved are also reasonable and in agreement with the modeling predictions of Lavvas et al. (2019) made for the temperature sub-Neptune GJ 1214 b. For the retrieval considering exo1 tholins, we constrain the particle radius in log unit to -1.65 (in m), the haze density in log unit to 11.22 (m-3), the mean pressure of the haze layer in log unit to -0.04 (in Pa), and the dP of the haze layer in log unit to 2.05 (in Pa), see Figure 7. The retrievals considering hazes also tend to increase the atmospheric temperature (see Figure 6) which along with the haze properties and the methane abundance affects the strength of the spectral features. Although this result reflects a degenerate behavior between different parameters influencing extinction, its physical meaning is reasonable as haze absorption in the UV-Visible tend to increase the atmospheric temperature of the upper atmosphere (Arney et al., 2016). This increased temperature driven by the haze heating properties could inhibit the formation of methane clouds and prevent the co-existence with the solid organic hazes.
@@ Acknowledgements.
This project has received funding from the European Research Council (ERC) under the ERC OxyPlanets projects (grant agreement No. 101053033). This project acknowledges funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 679030/WHIPLASH). T.D. acknowledges support by the 2024 NOMIS-ETH postdoc fellowship program. T.D. thanks the NOMIS foundation and ETH Zurich for funding via the 2024 Fellowship program. Some of the data presented in this paper were obtained from the Mikulski Archive for Space Telescopes (MAST) at the Space Telescope Science Institute. The specific observations analyzed can be accessed via doi:10.17909/3ds1-8z15 and doi:10.17909/rx29-yw62. We would like to thank the anonymous referee for his pertinent comments, which improved the presentation of our results.References
@@
- Arney et al. (2016) Arney, G. N., Domagal‑Goldman, S. D., et al. 2016, Astrobiology, 16, 873
- Bell et al. (2022) Bell, T., Ahrer, E.-M., et al. 2022, The Journal of Open Source Software, 7, 4503
@@
- Changeat et al. (2025) Changeat, Q., Bardet, D., et al. 2025, arXiv e-prints, arXiv:2505.18715
- Charnay et al. (2021) Charnay, B., Blain, D., et al. 2021, A&A, 646, A171
- Chubb et al. (2021) Chubb, K. L., Rocchetto, M., et al. 2021, A&A, 646, A21
- Drant et al. (2024) Drant, T., Garcia‑Caurel, E., et al. 2024, A&A, 682, A6
- Drant et al. (2025) Drant, T., Sciamma-O’Brien, E., et al. 2025, A&Ain review
- Gao et al. (2023) Gao, P., Piette, A. A. A., et al. 2023, The Astrophysical Journal, 951, 96
@@
- Hu et al. (2025) Hu, R., Bello-Arufe, A., et al. 2025, arXiv e-prints, arXiv:2507.12622
@@
- Khare et al. (1984) Khare, B. N., Sagan, C., et al. 1984, Icarus, 60, 127
@@
- Lavvas et al. (2019) Lavvas, P., Koskinen, T., et al. 2019, The Astrophysical Journal, 878, 118
- Luque et al. (2025) Luque, R., Piaulet-Ghorayeb, C., et al. 2025, arXiv e-prints, arXiv:2505.13407
- Madhusudhan et al. (2025) Madhusudhan, N., Constantinou, S., et al. 2025, ApJ, 983, L40
- Madhusudhan et al. (2023) Madhusudhan, N., Sarkar, S., et al. 2023, ApJ, 956, L13
- Schmidt et al. (2025) Schmidt, S. P., MacDonald, R. J., et al. 2025, arXiv e-prints, arXiv:2501.18477
- Sciamma-O’Brien et al. (2023) Sciamma-O’Brien, E., Roush, T. L., et al. 2023, The Planetary Science Journal, 4, 121
@@
- Trainer et al. (2006) Trainer, M., Pavlov, A., et al. 2006, PNAS, 103, 18035
- Welbanks et al. (2025) Welbanks, L., Nixon, M. C., et al. 2025, arXiv e-prints, arXiv:2504.21788
- Wogan et al. (2024) Wogan, N. F., Batalha, N. E., et al. 2024, The Astrophysical Journal Letters, 963, L7
@@ Appendix A Dataset offset
All these offsets are fixed to reduce the parameter space and accelerate aerosol retrievals. A more detailed investigation of their effects could be carried out in future work.
@@ Appendix B Retrieval setup
The non-equilibrium chemistry retrieval was carried out using FRECKLL coupled with TauREx 3, without aerosols. Kzz has been set to 1010. Its value, within common range of values, has been shown in Jaziri et al. (2025) to not significantly affect the retrieval. The priors are detailed in Table 3. Due to its high computational cost and limited stability, we present only one non-equilibrium retrieval, applied to the MIRI LRS data, which contain relatively few data points. This allows us to assess the consistency of the MIRI LRS observations under non-equilibrium conditions. We used 800 live points for all retrievals.
@@ Appendix C Inconsistency between NIRISS SOSS/NIRSpec G395H and MIRI LRS observations
A retrieval using all molecules on the NIRISS SOSS/NIRSpec G395H data yields a metallicity of 40.56, demonstrating that, in our case, free-chemistry retrievals tend to bias metallicities toward lower values compared to non-equilibrium models. Following Welbanks et al. (2025), this bias arises, at least in part, because we excluded from the retrieval certain species that may be abundant and influence the mean molecular weight but do not exhibit prominent spectral features, such as N2, in order to improve computational efficiency. This effect is also seen in our results when comparing retrievals with the full set of molecules to those with the reduced set (see metallicities in Tables 4 and 2). Nevertheless, the MIRI LRS retrievals, performed with the same setup as the NIRISS SOSS/NIRSpec G395H analysis and therefore directly comparable, tend to favor lower metallicity (from 40.56 to 4.42) to increase the scale height, allowing better fits to the larger observed features.
These results are inconsistent with those from the NIRISS SOSS/NIRSpec G395H analysis and the non-equilibrium chemical model, both of which favor higher metallicity and a high CH4 abundance, as seen in Figure 3. In contrast, the MIRI LRS retrievals indicate low metallicity and do not show a preference for CH4. However, as shown in Table 4, none of the tested models are statistically favored over a flat line. At best, they fall within the ”weak preference” range. H2CO is unlikely to be dominant. Its retrieved abundance of 10-5 exceeds the 3 upper limit derived by Jaziri et al. (2025). C2H2 and C2H4 could plausibly be present, especially considering high C/O ratio and high metallicity. However, the retrievals favor C2H2 over CH4, which is inconsistent with the NIRISS SOSS/NIRSpec G395H data. Our conclusions are consistent with previous analyses of MIRI LRS data using alternative reduction pipelines (Madhusudhan et al. 2025; Taylor 2025; Welbanks et al. 2025), which likewise report no significant spectral signatures. A possible exception is a weak indication of H2CO, though this is not statistically robust in a Bayesian framework and may simply reflect reduction-dependent systematics. In particular, features around 5-6 m, where H2CO absorption could occur, appear sensitive to the choice of data reduction, limiting the reliability of this tentative signal.